The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                    mcnews@astro.uiuc.edu
                         http://www.astro.uiuc.edu/mcnews/MCNews.html
                            http://www.astro.uni-bonn.de/~mcnews/

No. 23 August 3, 1998

Contents

News  1
Abstracts of 9 refereed papers  2
Abstract of 1 non-refereed paper 8

News

IAU Symposium 190 “New Views of the Magellanic Clouds”

The IAU Symposium 190 “New Views of the Magellanic Clouds” was held on the beautiful campus of the University of Victoria, Canada, on July 13–17, in honor of Dr. Sidney van den Bergh. 180 astronomers from 23 countries participated. 53 talks and 126 posters were presented covering the latest results from satellite and ground-based observations in all wavelength ranges as well as from theoretical models. A summary of the highlights from the meeting will be included in the next issue of the Magellanic Clouds Newsletter.

The manuscripts for the meeting proceedings are due on September 30. Instructions for the manuscript submission and other information about the meeting can be found in the web pages at

http://cadcwww.hia.nrc.ca/iau190/.

Many thanks to the Local Organizing Committee for their hospitality and for the perfect organization of a very enjoyable and successful meeting! We hope the next symposium on the Magellanic Clouds will be held in a not-too-distant future.

Eva Grebel & You-Hua Chu

Radio survey of X-ray sources available at CDS

FITS files of all the 28 deep radio images of fields in the Magellanic Clouds, from Fender, Southwell & Tzioumis ‘A radio survey of supersoft, transient and persistent X-ray sources in the Magellanic clouds’, MNRAS, 1998, 298, 692 are available for downloading at

                 ftp://cdsarc.u-strasbg.fr/pub/cats/J/MNRAS/298/692/

These include images of the SNR N49 (possibly associated with SGR 0525-66).

Robert P. Fender, Astronomical Institute ‘Anton Pannekoek’

Center for High-Energy Astrophysics, University of Amsterdam

The Netherlands

Abstracts of Refereed Papers

Evolutionary Models for the Magellanic Clouds: I. The Large Cloud

J.A. de Freitas Pacheco

Observatoire de la Côte d’Azur, BP4299, F-06304, Nice Cedex 4, France

An evolutionary model is developed for the Large Magellanic Cloud, taking into account constraints in the star formation history imposed by recent data on color-magnitude diagrams of field stars. The present model has a “closed” geometry, since the onset of a galactic wind in the early phases of the LMC is questioned. The code followed the evolution of several primary elements as O, Mg, Si, S, Ar, Ca and Fe, including the contribution of both type Ia and type II supernovae to the chemical enrichment. An initial mass function steeper than the solar neighborhood seems to be necessary to produce negative [O/Fe] ratios, but this conclusion depends on the adopted yields and on the adopted star formation history. The signature of the past enhanced star formation period 2-3 Gyr ago, is clearly seen in the behavior of the oxygen and sulphur evolution, representing a possible observational test for the proposed model.

Accepted by: The Astronomical Journal For preprints, contact pacheco@obs-nice.fr

The LMC Stellar Complexes in Luminosity Slices Star formation indicators

F. Maragoudaki^{1,2}, M. Kontizas^{1}, E. Kontizas^{2}, A. Dapergolas^{2} and D.H. Morgan^{3}

1 Section of Astrophysics, Astronomy & Mechanics, Dept. of Physics, University of Athens, GR-157 83 Athens, Greece 2 Astronomical Institute, National Observatory of Athens, P.O. Box 20048, GR-118 10 Athens, Greece 3 UK Schmidt Telescope Unit, Royal Observatory of Edinburgh, EH9 3HJ, UK

An approach towards the investigation of the star formation mechanisms in galaxies can be achieved through the search of stellar complexes and the determination of their properties. A method has been developed for the detection of stellar complexes and the derivation of their fundamental properties in the LMC.

Photographic plates taken with the 1.2m U.K. Schmidt Telescope have been digitized by the APM and SuperCosmos machines to produce homogeneous data for extended regions. Star counts have been performed for extended areas in selected luminosity slices and colours (U, HeII, R filters). Isodensity contours have been used to identify the various structures with enhanced stellar number density (3σ above the mean background density).

About 50 large stellar groupings have been revealed showing: 1) hierarchical structure, where the smallest are found within the large ones. 2) their size distribution has peaks at 250± 50pc (aggregates) and 600± 50pc (complexes), there are also a few cases with size 1200± 50pc (supercomplexes). 3) there is evidence that at the fainter magnitudes these structures are aligned to a general trend whereas at the most bright end (B1, O spectral types) they become more clumpy and symmetrical in shape. The relationship between the above complexes’ properties and the various scenarios of star formation are discussed.

Accepted by: Astronomy and Astrophysics (Letters)

For preprints, contact fmarag@titan.astro.noa.gr

An Empirical Isochrone of Very Massive Stars in R136A

Alex de Koter^{1,2}, Sara R. Heap^{3} and Ivan Hubeny^{3,4}

1 Advanced Computer Concepts, Code 681, Goddard Space Flight Center, Greenbelt MD 20771, USA 2 Astronomical Institute ‘Anton Pannekoek’, University of Amsterdam, Kruislaan 403, NL-1098 SJ Amsterdam, The Netherlands 3 Laboratory for Astronomy and Solar Physics, Code 681, Goddard Space Flight Center, Greenbelt MD 20771, USA 4 AURA, National Optical Astronomy Observatories, Code 681, Goddard Space Flight Center, Greenbelt MD 20771, USA

We report on a detailed spectroscopic study of twelve very massive and luminous stars (M ∼ > 35M_{⊙}) in the core of the compact cluster R136a, near the center of the 30 Doradus complex. The three brightest stars of the cluster, viz: R136a1, R136a2 and R136a3, have been investigated earlier by de Koter, Heap, & Hubeny (1997). Low-resolution spectra (< 200 km·s^{−1}) of the program stars were obtained with the GHRS and FOS spectrographs on the Hubble Space Telescope. These instruments covered the spectral range from 1200 to 1750 Å and from 3200 to 6700 Å respectively. Fundamental stellar parameters were obtained by fitting the observations by model spectra calculated with the unified isa-wind code of de Koter et al. supplemented by synthetic data calculated using the program tlusty.

We find that the stars are almost exclusively of spectral type O3. They occupy only a relatively narrow range in effective temperatures between 40 and 46 kK. The reason for these similar T_{eff}’s is that the isochrone of these very massive stars, which we determined to be at ∼ 2 Myr, runs almost vertically in the HR-diagram. We present a quantitative method of determining the effective temperature of O3-type stars based on the strength of the O v λ1371 line.

Present-day evolutionary calculations by Meynet et al. (1994) imply that the program stars have initial masses in the range of M_{i} ∼ 37 to 76 M_{⊙}. The observed mass-loss rates are up to three (two) times higher than is assumed in these evolution tracks when adopting a metallicity Z = 0.004 (0.008) for the LMC. The high observed mass-loss rates imply that already at an age of ∼ 2 Myr the most luminous of our program stars will have lost a significant fraction of their respective initial masses.

For the least luminous stars investigated in this paper, the observed mass loss agrees with the prediction by the theory of radiation-driven winds (Kudritzki et al. 1989). However, for increasing luminosity the observed mass loss becomes larger, reaching up to three to four times what is expected from the theory. Such an increasing discrepancy fits in with the results of de Koter et al. (1997) where an observed over predicted mass loss ratio of up to eight was reported for the brightest members of the R136a cluster, for which M_{i} ∼ 100 M_{⊙} was found. The failure of the theory is also present when one compares observed over predicted wind momentum as a function of wind performance number. This strongly indicates that the shortcoming of the present state of the theory is connected to the neglect of effects of multiple photon momentum transfer.

Accepted by: The Astrophysical Journal

For preprints, contact dekoter@astro.uva.nl
or by anonymous ftp at helios.astro.uva.nl, cd pub/alex/r136aempiso.ps

A slope variation in the Period-Luminosity relation for short period SMC Cepheids

F. Bauer^{1,2}, C. Afonso^{2}, J.N. Albert^{3}, J. Andersen^{4}, R. Ansari^{3}, E. Aubourg^{2}, P. Bareyre^{1,2}, J.P. Beaulieu^{5}, A. Bouquet^{1}, S. Char^{6}, X. Charlot^{2}, F. Couchot^{3}, C. Coutures^{2}, F. Derue^{3}, R. Ferlet^{5}, C. Gaucherel^{2}, J.F. Glicenstein^{2}, B. Goldman^{2,7,8}, A. Gould^{9}, D. Graff^{2,10}, M. Gros^{2}, J. Haissinski^{3}, J.C. Hamilton^{1}, D. Hardin^{2}, J. de Kat^{2}, T. Lasserre^{2}, E. Lesquoy^{2}, C. Loup^{5}, C. Magneville^{2}, B. Mansoux^{3}, J.B. Marquette^{5}, E. Maurice^{1}1, A. Milsztajn^{2}, M. Moniez^{3}, N. Palanque-Delabrouille^{2}, O. Perdereau^{3}, L. Prevot^{1}1, C. Renault^{2},

N. Regnault^{3}, J. Rich^{2}, M. Spiro^{2}, A. Vidal–Madjar^{5}, L. Vigroux^{2}, S. Zylberajch^{2}

The EROS collaboration

1 Collège de France, PCC, IN2P3 CNRS, 11 place Marcelin Berthelot, 75231 Paris Cedex, France 2 CEA, DSM, DAPNIA, Centre d’Études de Saclay, 91191 Gif-sur-Yvette, Cedex, France 3 Laboratoire de l’Accélérateur Linéaire, IN2P3 CNRS et Université Paris-Sud, BP 34 91898 Orsay Cedex, France 4 Astronomical Observatory, Copenhagen University, Juliane Maries Vej 30, 2100 Copenhagen, Denmark 5 Institut d’Astrophysique de Paris, INSU CNRS, 98 bis Boulevard Arago, 75014 Paris, France 6 Universidad de la Serena, Facultad de Ciencias, Departamento de Fisica, Casilla 554, La Serena, Chile 7 Dept. Astronomía, Universidad de Chile, Casilla 36-D, Santiago,Chile 8 European Southern Observatory, Casilla 19001, Santiago 19, Chile 9 Department of Astronomy, Ohio State University, Columbus, OH 43210, USA 10 Physics Department, Ohio State University, Columbus, OH 43210, USA 11 Observatoire de Marseille, 2 place Le Verrier, 13248 Marseille Cedex 04, France

We present the Period–Luminosity relations from 290 Cepheids towards the LMC and 590 Cepheids towards the SMC. The two data sets were obtained using the two wide field CCD cameras of the EROS 2 microlensing survey. We observe a significant slope change of the period–luminosity relation for the SMC fundamental mode Cepheids with periods shorter than 2 days. Many possible experimental biases have been investigated, but none can account for this effect. We discuss three possible explanations of this slope change.

Submitted to: Astronomy & Astrophysics For preprints, contact bauer@cdf.in2p3.fr Also available from: astro-ph/9807094

Optical Gravitational Lensing Experiment. Population Effects on the Mean Brightness of the Red Clump Stars

A. Udalski

Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland

We present an empirical test indicating that the mean I-band magnitude of the red clump stars, used as the standard candle in the recent distance determinations to the Magellanic Clouds and other objects, is age independent for intermediate age (2 − 10 Gyr) stars. Fifteen star clusters of age ≈ 1.5 − 12 Gyr from the LMC and SMC (ESO121SC03, SL663, NGC 2155, NGC 2121, SL388, SL862, NGC 121, L1, Kron 3, NGC 416, L113, NGC 339, L11, NGC 419, NGC 411) were observed and their color-magnitude diagrams are presented. The mean I-band brightness of the red clump in these clusters is constant and its mean extinction-free magnitude is: I_{0} = 17.88 ± 0.05 mag and I_{0} = 18.31±0.07 mag at the mean metallicity of −0.8 dex and −1.2 dex for the LMC and SMC clusters, respectively. For older objects (> 10 Gyr) the brightness of the red clump, which converts into the red part of the horizontal branch, fades by about 0.3 − 0.4 mag, setting an important limitation on the red clump stars method of distance determination.

The red clump distance moduli to the Magellanic Clouds from the new independent data set are: m − M = 18.18 ± 0.06 mag and m − M = 18.65 ± 0.08 mag for the LMC and SMC, respectively, in very good agreement with previous determinations.

Weak dependence of the mean I-band brightness of the red clump on metallicity and its independence of age for intermediate age population (2 − 10 Gyr) of stars as well as the most precise calibration as compared to other standard candle candidates makes the red clump stars method one of the most accurate steps in the distance scale ladder.

Submitted to: Acta Astronomica

For preprints, contact udalski@sirius.astrouw.edu.pl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9807095

On the precession of accretion discs in X-ray binaries

J. Larwood^{1,2}

1 Isaac Newton Institute for Mathematical Sciences, 20 Clarkson Road, Cambridge CB3 0EH. United Kingdom 2 Astronomy Unit, School of Math. Sciences, Queen Mary & Westfield College, Mile End Road, London E1 4NS, UK

In this letter recent results on the nodal precession of accretion discs in close binaries are applied to the discs in some X-ray binary systems. The ratio between the tidally forced precession period and the binary orbital period is given, as well as the condition required for the rigid precession of gaseous Keplerian discs. Hence the minimum precessional period that may be supported by a fluid Keplerian disc is determined. It is concluded that near rigid body precession of tilted accretion discs can occur and generally reproduce observationally inferred precession periods, for reasonable system parameters. In particular long periods in SS433, Her X-1, LMC X-4 and SMC X-1 can be fit by the tidal model. It is also found that the precession period that has been tentatively put forward for Cyg X-2 cannot be accommodated by a tidally precessing disc model for any realistic choice of system parameters.

Accepted by: Monthly Notices of The Royal Astronomical Society

For preprints, contact larwood@mpia-hd.mpg.de
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9806348

EROS 2 intensive observation of the caustic crossing of microlensing event MACHO SMC-98-1

Afonso et al. (EROS collaboration)

We report on intensive photometric monitoring on 18 June 1998 of MACHO SMC-98-1, a binarylens microlensing event seen toward the Small Magellanic Cloud (SMC). The observations cover 5.3 hours (UT 5:17 – 10:37), and show a sharp drop of 1.8 mag during the first 1.8 hours, followed by an abrupt flattening at UT 7:08 ± 0:02. We interpret the kink at 7:08 as the end of the second caustic crossing (when the source first moved completely outside the caustic). These results indicate that μ sin ϕ< ∼ 1.5 km s^{−1} kpc^{−1} at the 2σ level, where μ is the proper motion of the lens (relative to the line of sight to the source), and ϕ is the unknown (and so random) angle of the caustic crossing. Hence, the lens¸ probably does not lie in either the Galactic halo or disk and so is most likely in the SMC itself. Our data can be combined with those of other groups to give more precise constraints on the proper motion (and hence the nature) of the lens.

Submitted to: Astronomy and Astrophysics

For preprints, contact Nathalie.Delabrouille@cea.fr
Also available from the URL ftp://ftp.lal.in2p3.fr/pub/Eros/Papiers/caustic.ps.gz

The Relative Lens-Source Proper Motion in MACHO 98-SMC-1

M.D. Albrow^{1}, J.-P. Beaulieu^{2}, J.A.R. Caldwell^{3}, D.L. DePoy^{4}, M. Dominik^{2}, B.S. Gaudi^{4}, A. Gould^{4}, J. Greenhill^{5}, K. Hill^{5}, S. Kane^{5,6}, R. Martin^{7}, J. Menzies^{3}, R.M. Naber^{2}, J.-W. Pel^{2}, K.R. Pollard^{1}, P.D. Sackett^{2}, K.C. Sahu^{6}, P. Vermaak^{3},

R. Watson^{5}, A. Williams^{7}, (The PLANET Collaboration) and R.W. Pogge^{4}

1 Univ. of Canterbury, Dept. of Physics & Astronomy, Private Bag 4800, Christchurch, New Zealand 2 Kapteyn Astronomical Institute, Postbus 800, 9700 AV Groningen, The Netherlands 3 South African Astronomical Observatory, P.O. Box 9, Observatory 7935, South Africa 4 Ohio State University, Department of Astronomy, Columbus, OH 43210, USA 5 Univ. of Tasmania, Physics Dept., G.P.O. 252C, Hobart, Tasmania 7001, Australia 6 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 7 Perth Observatory, Walnut Road, Bickley, Perth 6076, Australia

We present photometric and spectroscopic data for the second microlensing event seen toward the Small Magellanic Cloud (SMC), MACHO-98-SMC-001. The lens is a binary. We resolve the caustic crossing and find that the source took 2Δt = 8.5 hours to transit the caustic. We measure the source temperature T_{eff} = 8000 K both spectroscopically and from the color (V − I)_{0} ∼ 0.22. We find two acceptable binary-lens models. In the first, the source crosses the caustic at ϕ = 43^{∘}.2 and the unmagnified source magnitude is I_{s} = 22.15. The angle implies that the lens crosses the source radius in time t_{∗} = Δt sin ϕ = 2.92 hours. The magnitude (together with the temperature) implies that the angular radius of the source is θ_{∗} = 0.089 μas. Hence, the proper motion is μ = θ_{∗}/t_{∗} = 1.26 km s^{−1} kpc^{−1}. For the second solution, the corresponding parameters are ϕ = 30^{∘}.6, I_{s} = 21.81, t_{∗} = 2.15 hours, θ_{∗} = 0.104 μas, μ = θ_{∗}/t_{∗} = 2.00 km s^{−1} kpc^{−1}. Both proper-motion estimates are slower than 99.5% of the proper motions expected for halo lenses. Both are consistent with an ordinary binary lens moving at ∼ 75–120 km s^{−1} within the SMC itself. We conclude that the lens is most likely in the SMC proper.

Submitted to: Astrophysical Journal Letters

For preprints, contact gould@astronomy.ohio-state.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9807086

Discovery and Characterization of a Caustic Crossing Microlensing Event in the SMC

C. Alcock, R.A. Allsman, D. Alves, T.S. Axelrod, A.C. Becker, D.P. Bennett, K.H. Cook, A.J. Drake, K.C. Freeman, K. Griest, L.J. King, M.J. Lehner, S.L. Marshall, D. Minniti, B.A. Peterson, M.R. Pratt, P.J. Quinn, S.H. Rhie,

A.W. Rodgers, P.B. Stetson, C.W. Stubbs, W. Sutherland, A. Tomaney, T. Vandehei

We present photometric observations and analysis of the second microlensing event detected towards the Small Magellanic Cloud (SMC), MACHO Alert 98-SMC-1. This event was detected early enough to allow intensive observation of the lightcurve. These observations revealed 98-SMC-1 to be the first caustic crossing, binary microlensing event towards the Magellanic Clouds to be discovered in progress.

Frequent coverage of the evolving lightcurve allowed an accurate prediction for the date of the source crossing out of the lens caustic structure. The caustic crossing temporal width, along with the angular size of the source star, measures the proper motion of the lens with respect to the source, and thus allows an estimate of the location of the lens. Lenses located in the Galactic halo would have a velocity projected to the SMC of v̂ ∼ 1500 km s^{−1}, while an SMC lens would typically have v̂ ∼ 60 km s^{−1}. The event lightcurve allows us to obtain a unique fit to the parameters of the binary lens, and to estimate the proper motion of the lensing system.

We have performed a joint fit to the MACHO/GMAN data presented here, including recent EROS data of this event. These joint data are sufficient to constrain the time t_{∗} for the lens to move an angle equal to the source angular radius; t_{∗} = 0.116 ± 0.010 days. We estimate a radius for the lensed source of R_{∗} = 1.4 ± 0.1R_{⊙} from its unblended color and magnitude. This yields a projected velocity of v̂ = 84 ± 9 km s^{−1}. Only 0.15 % of halo lenses would be expected to have a v̂ value at least as small as this, while 31% of SMC lenses would be expected to have v̂ as large as this. This implies that the lensing system is more likely to reside in the SMC than in the Galactic halo. Similar observations of future Magellanic Cloud microlensing events will help to determine the contribution of Machos to the Galaxy’s dark halo.

Submitted to: The Astrophysical Journal

For preprints, contact becker@astro.washington.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9807163

Abstracts of Non-Refereed Papers

Superbubbles in the Magellanic Clouds

M. S. Oey^{1}

1 Institute of Astronomy, Madingley Road, Cambridge CB3 0HA, U.K.

Superbubbles that result from the stellar winds and supernovae of OB associations probably play a fundamental role in the structure and energetics of the ISM in star-forming galaxies. Their influence may also dominate the relationship between the different interstellar gas phases. How do superbubbles form and evolve? How do they affect the local and global ISM? The Magellanic Clouds provide a superior opportunity to study this shell-forming activity, since both stellar content and gaseous structure can be examined in detail. Here, the results of recent studies of superbubbles in the Magellanic Clouds are reviewed.

Review paper to appear in New Views of the Magellanic Clouds, IAU Symp. 190 (July 12-17, 1998), eds. Y.-H. Chu, N. Suntzeff, J. Hesser, & D. Bohlender.

Preprints from oey@ast.cam.ac.uk
Also available from the URL http://www.ast.cam.ac.uk/~oey/oeypubs.html
or http://xxx.lanl.gov/abs/astro-ph/9807271